### v1

# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

Abstract
Networks of physical units can vary from a stationary set of spatially-embedded links to a collection of mobile agents that undergo transient social interactions. In living cells, mitochondria form architectures that span across these regimes, transitioning between fragmented, partly connected, and highly fused structures depending on cell type and state. Diffusive transport of biomolecular components through these networks helps to homogenize the mitochondrial population. Here we address the connection between dynamic network architecture and the rate of diffusive mixing through simulations and analytic models that incorporate fusion, fission, and rearrangement. We find that the material delivered from a source to the rest of the network depends on the network dimensionality and a balance of competing timescales for encounter, fusion, and diffusive dispersion. These results provide a quantitative basis for predicting the homogenization of proteins, lipids, ions, or genetic material through the mitochondrial population. The general principles identified in this work capture diffusive spreading through both social and physical networks, unifying a continuum of spatial network architectures.
Introduction
Diffusive transport through networks has been studied in a variety of contexts, including disease spread in epidemiology [1, 2], innovations in social networks [3], communication among insects [4, 5], nutrient flows in fungal networks [6], diffusion of signals and nutrients in the brain extracellular space [7, 8], and oil recovery in porous rock [9]. These systems can be separated into two broad classes: one where the topology of network connections is stationary and limited to nearest neighbors, and another where network nodes are mobile and promiscuous, interacting with many different partners over time.
Physical networks form an important category of stationary, spatially-constrained network structures [10]. In such networks, edges represent objects subject to physical limitations (space-filling, steric repulsion, etc.) [11, 12] connecting degree-1 tips and degree-3 junctions, with higher degree nodes exceedingly rare [13, 14]. Examples include fungal mycelia [15], porous rocks [16], and neuronal synaptic networks [17]. Transport behaviors on stationary networks can be described by the graph Laplacian, whose eigenvalues govern the spreading timescales [18, 19, 20]. In some percolation problems, physical networks become dynamic as edges are allowed to flicker between an active and inactive state [21, 22].
A distinct set of approaches considers signal transmission in ‘social’ networks of transiently-interacting mobile units. In such networks, the mixing behavior is defined by distributions of contact durations and inter-contact times [4, 23], which may be constrained by spatial embedding [2, 1]. Notably, many previously studied social network systems focus on the spreading of a non-diluting signal (infection, information, etc.) [4, 2, 3], in contrast to dispersion of mass-conserving physical material [24, 22, 25].
Despite the distinct modeling approaches employed, physical and social networks lie on a continuum of temporal network structures [23] with varying timescales of topological rearrangement. The intracellular environment exemplifies this spectrum. The endoplasmic reticulum forms a highly-looped physical lattice of membrane-bound tubules that enables diffusive transport of proteins [25] and ions [26] across the entire cell. At the opposite extreme, the population of endocytic vesicles constitutes a social network of discrete transiently-interacting compartments [27, 28].
Mitochondria form another intracellular network with striking structural variability. Mitochondrial networks transition between fragmented and hyperfused architectures in response to disease state [29, 30, 31, 32, 33, 34], metabolic conditions [35, 36, 37], calcium signalling and apoptosis [38, 39], cell division [40, 41] and cell type [42, 43, 44, 45]. These structural changes are thought to support important functions such as genetic complementation [46, 47], dilution of harmful reactive oxygen species (ROS) [48], power cabling [49, 50], quality control [51, 52, 53, 54], and modulation of mitochondrial heterogeneity across the population [55, 56]. These cellular functions rely on the connectivity-dependent diffusive spread of proteins, lipids, and/or ions through the underlying mitochondrial network [57, 22].
To elucidate how mitochondrial structure and dynamics govern diffusive mixing, we introduce a quantitative framework for transport on spatially-embedded dynamic networks. We show that network connectivity governs spreading behavior by modulating several competing timescales: the cluster filling rate, encounter rate, and fusion rate, as well as the decay rate of the spreading material. In particular, we demonstrate a transition from rapid but low-dimensional spread in well-connected networks to slower three-dimensional spreading through an interacting population. Due to the different dimensionalities involved, particles with distinct diffusivities and networks with distinct structures can exhibit qualitatively different scaling behavior for dispersion from a source. The network architectures considered here are motivated by observations of three-dimensional mammalian mitochondria, whose structural diversity exemplifies a broad range of spatial networks.
Figure 1: Dynamic network simulation framework. (a) A network is formed by interacting units which diffuse through 3D space with diffusivity $D_{1}$. Nearby units can undergo tip-tip and tip-side fusion with rates $k_{u1},k_{u2}$ while connected nodes undergo fission at rate $k_{f},\frac{3}{2}k_{f}$, respectively. Material spreads along connected units with diffusivity $D_{p}$ and decays over time with rate $k_{d}$. Red indicates the concentration of material in a given unit. (b) Snapshots of spreading on simulated networks at steady state. (i) Fragmented network, ($k_{u1}/k_{f}=30$) (ii) Network near percolation transition, with kinetic parameters appropriate to mammalian mitochondrial networks [58] ($k_{u1}/k_{f}=1000$). (iii) Hyperfused network ($k_{u1}/k_{f}=3000$). The ratio of tip-tip and tip-side fusion is set to $k_{u1}=3k_{u2}$, and particle diffusivity is $D_{p}=4800$ throughout.
Model description
We seek to quantify the rate of material dispersion through dynamic spatial networks ranging from the social to the physical regime. These limiting architectures, as well as a continuum of intermediate network connectivities, are encompassed by a simulation framework [58], illustrated in Fig. 1, consisting of $N_{0}$ interacting spherocylindrical units of length $\ell_{0}$, which move diffusively through a bounded spherical domain of radius $R$ with unit diffusivity $D_{1}$. Upon close encounters, pairs of units may fuse into larger clusters with rate constants $k_{u1}$ (for tip-tip fusions) or $k_{u2}$ (for tip-side fusions). Fission breaks connections at the nodes, with fission rate proportional to the number of attached edges ($k_{f}$ at degree-2 nodes and $1.5k_{f}$ at degree-3 nodes), in keeping with prior models [14]. At steady state, the system exhibits a characteristic mean cluster size $\langle n\rangle$ set by the balance of fusion and fission, and an associated cluster diffusivity $D_{n}$. Details of this structural model and its relevant parameterization for mitochondrial networks in different cell types are described in Ref. [58]; here we focus on the spreading of material throughout the resulting dynamic networks.
We select at random a single unit, which serves as a particle source with fixed constant concentration $h_{0}=1/\ell_{0}$. Such a fixed source could represent, for instance, a mitochondrion that buffers calcium at a contact site with the endoplasmic reticulum [59], or one that produces a regulated amount of mitochondrially encoded The concentration field is propagated between connected units on the network with diffusivity $D_{p}$, using a finite volume approach [60] discretized at the level of individual units. Concentrations decay with a constant rate $k_{d}$ (representing removal from the network). The decay rate serves to set a relevant timescale, with the resulting calculations closely related to the question of how much material spreads from the source over a certain time. We quantify dispersion by computing the total amount of material ($S$) in the network at steady-state, excluding the source unit. Example steady-state snapshots are shown in Fig. 1b and Supplemental Video 1.
Given the complex internal structure of mitochondria and the variety of functionally relevant biomolecules within them, particle diffusivities can vary broadly. For example, proteins in the inner mitochondrial membrane can become trapped in the extensive folds of cristae [61]. The cristae also serve as diffusive barriers to hinder transport of solutes in the matrix [62] and intermembrane space [63]. Our simplified modeling approach coarse-grains these complications into a single effective diffusivity $D_{p}$ representing the motion of a particle along the mitochondrial tubule axis. Reported effective diffusivities vary over orders of magnitude, from $\sim 0.004\mu\text{m}^{2}/\text{s}$ for ATP synthase components  [64], to $\sim 20\mu\text{m}^{2}/\text{s}$ for mitochondrial matrix proteins [65], with diffusivity of ions presumed even higher [66]. We consider a range of diffusivities $D_{p}=0.4-40\mu\text{m}^{2}/\text{s}$ in our simulations. The diffusivity of individual mitochondria is estimated to be much slower, at $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$ [42, 67].
The results below are reported in dimensionless units, relative to the length scale $2\ell_{0}$ and the timescale $1/k_{f}$. A reasonable estimate sets the mitochondrial unit length to $\ell_{0}=0.5\mu$m and the fission timescale to $1/k_{f}\approx 2\text{ min}$ [58], consistent with experimental measurements of overall fission rate [68, 42]. The corresponding dimensionless diffusivities are $D_{p}=48-4800$ for the material, and $D_{1}=0.5$ for individual mitochondria.
Results
Spreading on the network is governed by the interplay of several key timescales: the decay time $\tau_{\text{d}}=1/k_{d}$, the cluster filling time $\tau_{\text{c}}$ for particles to diffusively explore a typical-sized connected component (cluster), and the waiting time between interactions $\tau_{\text{int}}$ (the time to encounter and fuse with a new cluster). The fission time $\tau_{f}=1/k_{f}$ sets the typical duration of transient interactions between distinct clusters. The material exhibits different dynamics depending on the comparative values of these four timescales. We begin by considering several limiting regimes where different timescales dominate.
Highly connected regime: spreading through stationary networks
In the limit where $\tau_{d}$ is the shortest timescale, material spreads primarily within a single connected component. This regime is relevant for highly fused networks with large cluster sizes. Such networks also tend to be relatively static in their topology, with any fissions that occur rapidly followed by re-fusion with the same neighboring unit [58]. For static networks, the steady-state distribution of material can be found by solving the diffusion equation on each one-dimensional edge, while matching boundary conditions at each node (details in SI Appendix). This method provides an exact solution for each individual network structure. However, to gain insight on the relevant scaling regimes, we turn to a mean-field continuum approach.
Figure 2: Material spreading on static networks of different dimension $d$. Two different sets of networks are considered: near-linear structures (blue) with $k_{u2}=0.01k_{u1}$ and highly-branched structures (green) with $k_{u2}=10k_{u1}$. The simulations are run to steady state, and the network structures are then frozen. (a) Calculation of the network dimension for a single instance of each network type (shown in inset). The number of nodes within a given graph distance is plotted against the graph distance on log-log axes, with the slope giving the dimension, $d$. Averaging over 21 snapshots from 3 independent simulations yields effective dimensionalities $d=1.2,d=1.9$, respectively. (b) Total material delivered from a source unit on the static networks is plotted as a function of the diffusive lengthscale, $\lambda_{p}=\sqrt{D_{p}/k_{d}}$. Solid curves show exact solution (details in SI Appendix), averaged over replicate snapshots. Dashed-circle curves show approximation with the linear motif (inset i), applicable at short $\lambda$. Dashed-square curves show continuum solution on a fractal domain (inset ii), matching well at long $\lambda$.
Depending on the relative rates of tip-tip versus tip-side fusion, network architectures can range from near-linear snake-like structures to highly branched compact morphologies, with the latter allowing for more rapid spreading of material through the network (Fig. 2). These structural features can be described by an effective fractal dimension $d$, which sets the scaling relationship between the number of units and the network ‘size’ in terms of graph distance. Specifically, we count the number of nodes within a given graph distance from a starting node, averaged over all possible starting nodes [69]. The power-law scaling exponent of this curve defines the dimension $d$ (Fig. 2i, inset) and therefore the cluster filling timescale, $\tau_{c}=(\left<n\right>^{1/d}\ell_{0})^{2}/D_{p}$. Although the simulated networks are embedded in a three-dimensional space, the limited junction degree gives rise to networks that are of lower dimension ($1\leq d\leq 2$). This feature has been previously observed in a variety of physical networks, from plant roots to ant tunnels, composed of tube-like objects connected at junctions [13].
[…]
For the regime where material spreads over a more extensive network structure ($\lambda_{p}$ much greater than segment length), we consider diffusion through a $d$-dimensional medium. Here we approximate the spreading as spherically symmetric, with a $d-$sphere of radius $a_{n}$ (enclosing the source edge) maintained at fixed concentration, and a reflecting $d-$sphere of radius $R_{n}$ representing the outer boundary of the cluster. We set $R_{n}$ such that the average distance between two points in the $d$-dimensional domain is equal to the average graph distance between network nodes (details in SI Appendix). Assuming that the $n\right>$ network units are uniformly distributed within the continuum sphere, the radius $a_{n}$ is set proportionately to allow for 1 source unit within the inner sphere: $1/a_{n}^{d}=(\langle n\rangle-1)/(R_{n}^{d}-a_{n}^{d})$.
We can write down the diffusion equation in $d$-space, $\left(\frac{1}{d}D_{p}\right)\frac{1}{r^{d-1}}\frac{\partial}{\partial r}\left(r^{d-1}\frac{\partial c(r)}{\partial r}\right)-k_{d}c(r)=0$, where the effective particle diffusivity is scaled by the dimension $d$. This scaling is applied because, when diffusing along an edge, the particle only moves in one dimension at each instance in time rather than simultaneously in all $d$ dimensions. The $d$-space diffusion equation has solutions in terms of modified Bessel functions $I_{\nu}(x),K_{\nu}(x)$ [70], where $\nu=1-d/2$ and $x=r\sqrt{d}/\lambda_{p}$. Taking a fixed concentration boundary at $x_{1}=a_{n}\sqrt{d}/\lambda_{p}$ and a reflecting boundary at $x_{2}=R_{n}\sqrt{d}/\lambda_{p}$ we find the total material delivered (details in SI Appendix):
$\begin{gathered}S_{d}=\frac{d}{x_{1}}\frac{\left[-I_{\nu-1}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu-1}(x_{1})\right]}{\left[I_{\nu}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu}(x_{1})\right]}\\ \end{gathered}$ (1)
Fig. 2 shows the correspondence between diffusion through a fractal continuum (dashed-square curves) and the exact solution (solid curves) for networks with different connectivities. The linear motif solution $S_{1}$ (dashed-circle curves) is a good estimate for small diffusive lengths, while the continuum solution $S_{d}$ (dashed-square curves) is a better approximation for large $\lambda_{p}$. When the diffusive lengthscale is large enough ($\lambda_{p}>R_{n}$), the entire connected component saturates. Overall, material delivery is boosted by increased network branching [22], which increases both the network dimension and the network density (lower $R_{n}$ for the same number of network units).
Fragmented network regime: spreading through transient interactions
We next consider the regime of highly fragmented ‘social’ networks, where clusters are sufficiently small (and particle diffusivity sufficiently fast) that concentrations are fully equilibrated within each cluster: $\tau_{c}\ll\left\{\tau_{\text{d}},\tau_{\text{int}}\right\}$. In this limit, we can approximate the network as a system of $N=N_{0}/\left<n\right>$ identical spherical units with effective radius $a$, each representing a cluster of uniform concentration. These effective units are capable of undergoing transient fusions whenever they are within contact radius $b$, with rate constant $k_{u}$. Each fusion equilibrates the particle concentration in the two units involved, and is instantaneously followed by fission so that no larger structures are formed.
Results from an explicit simulation of this simplified model are shown in Fig. 3 (dots). A mean-field analytic approximation can be found by fixing the source unit in the center of a domain, assigning a diffusivity of $D=2D_{n}$ to the remaining units and solving for the spatial concentration field $h(r,t)$, which defines the mean concentration per network unit located at distance $r$ from the origin. The average of this field within the narrow contact zone is defined as $h_{c}(t)$. The time evolution of the concentration fields can be expressed as follows (details in SI Appendix):
$\frac{dh(r,t)}{dt}$ $=D\nabla^{2}h(r,t)-k_{d}h(r,t),\;\;\text{for}\ b<r<R$ (2a)
$\frac{dh_{c}(t)}{dt}$ $=k_{u}(h_{0}\ell_{0}-h_{c}(t))-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (2b)
where $k_{u}(h_{0}\ell_{0}-h_{c}(t))$ represents the injection of new material into the system via fusion with the source, $I=-4\pi b^{2}\rho D\left.\frac{dh(r,t)}{dr}\right|_{b}$ is the current of material leaving the contact zone, $\rho=(N_{0}-\langle n\rangle)/V$ is the density of network units, $V=4/3\pi(R^{3}-a^{3})$ is the domain volume, and $v_{c}=4/3\pi(b^{3}-a^{3})$ is the contact zone volume. Interactions between non-source units do not alter the mean-field concentrations, and the overall material in the system can only increase upon encounters with the source.
Figure 3: Material delivered from a source unit in a system of fragmented clusters. Solid lines show mean-field solutions (see Eq. 3) with unit cluster size $\langle n\rangle=1$, for different fusion rates $k_{u}$. Colored dots show explicit simulation results for a simplified system of interacting spheres with uniform size (inset). Dashed black lines show the limiting behavior for fusion-limited and diffusion-limited regimes.
We compute the steady-state solution of Eq. 2 and integrate $h(r)$ over the space to find the total amount of material ($S$) in the system, excluding the source unit (details in SI Appendix):
$S=$ $(\langle n\rangle-1)+\rho h_{c}(v_{c}+z),$ (3a)
$h_{c}=$ $\frac{1}{1+k_{d}/k_{u}+k_{d}z/k_{u}v_{c}},$ (3b)
$\frac{z}{4\pi b\lambda}=$ $\frac{(\lambda-b)(R+\lambda)+(\lambda+b)(R-\lambda)e^{2(R-b)/\lambda}}{(R+\lambda)+(R-\lambda)e^{2(R-b)/\lambda}},$ (3c)
where $\lambda=\sqrt{D/k_{d}}$ is the diffusive lengthscale for spreading through the population of clusters. Here, the first term in $S$ represents material in the cluster containing the source unit, the term $\rho h_{c}v_{c}$ is the total material within the contact layer (defined by a balance of fusion, decay, and diffusive escape), and $z$ describes the additional volume over which the material has spread beyond the contact zone.
The total material in the system exhibits distinct scaling behaviors (Fig. 3, dashed) depending on the relative timescales for decay, encounter ($\tau_{\text{enc}}=(4\pi Db\rho)^{-1}$), and fusion ($\tau_{u}=(k_{u}\rho v_{c})^{-1}$). In keeping with standard results for diffusion to a partially reactive target [71], the overall interaction time is given by the sum of waiting times for the two-step process: $\tau_{\text{int}}=\tau_{\text{enc}}+\tau_{u}$. In the limit of fast decay ($\tau_{d}\ll\tau_{\text{int}}$), the steady-state material in a social network of unit size clusters approaches $S\rightarrow\rho v_{c}h_{c}=\rho v_{c}k_{u}/k_{d}$. In this limit, the total amount delivered is set by the number of units in the contact zone ($\rho v_{c}$) and how often they fuse with the source during the decay time ($k_{u}/k_{d}$), with no spatial spread.
Another limit arises when diffusion of clusters is fast relative to both decay and fusion ($\tau_{\text{enc}}\ll\tau_{d},\tau_{u}$). This yields the same scaling for the total network content $S\rightarrow\rho v_{c}k_{u}/k_{d}$, with $h_{c}\rightarrow k_{u}v_{c}/4\pi Db$. The amount of material in the contact layer ($h_{c}$) is set by a balance between injection through fusion (at rate $k_{u}\rho v_{c}$) and escape through the diffusive arrival of fresh units that dilute the local concentration ($4\pi Db\rho$). The additional volume $z\rightarrow 4\pi Db/k_{d}$ over which the material spreads depends on the balance between diffusive encounters and decay. Overall, the total amount of material in the network is fusion-limited as the rapid encounters quickly homogenize the individual units within the diffusive range of the source.
In a third limit, fusion of clusters is fast and decay is slow relative to the timescale of diffusive encounter. ($\tau_{u}\ll\tau_{\text{enc}}\ll\tau_{d}$). For this regime, the total material content is given by $S\rightarrow\rho zh_{c}\rightarrow 4\pi Db\rho/k_{d}$. This is a diffusion-limited regime, where clusters arriving at the contact region fuse with the source nearly instantaneously ($h_{c}\approx h_{0}\ell_{0}=1$), and spreading is determined by the balance of arrival and decay rates.
Fig. 3 shows the full solution of the mean-field model (Eq. 3) for spreading on a fragmented network, with variable fusion rate $k_{u}$ and a unit cluster size $n\right>=1$. The analytic calculations accurately reproduce simulations with diffusive spheres that exchange material via transient fusion (dots in Fig. 3, details in SI Appendix).
Figure 4: Spreading on simulated networks exhibits a transition between the static network and social network regime. The total material is plotted as a function of the decay time for different fusion rate constants $k_{u1}$ (solid lines). Approximation of the network as a fractal continuum captures the behavior in the large cluster, fast decay limit, while the mean-field social network approximation describes the small cluster, slow decay limit. The combined solution (Eq. 4, dashed curves) encompasses the full range of decay timescales and particle diffusivities. The transition between the two approximations occurs when the cluster containing the source unit is filled ($S=\langle n\rangle-1$), as marked by colored arrows in (a). Insets show example simulation snapshots for the parameters indicated by the gray lines. Parameters $N_{0}=250,\ell_{0}=0.5,R=5,k_{u2}=\frac{1}{3}k_{u1},k_{f}=1$ are used in both (a) and (b), with the particle diffusivity reduced from $D_{p}=4800$ in (a) to $D_{p}=48$ in (b).
Dynamic networks with large interacting clusters
The model for spreading through a social network of interacting fragments can be expanded to approximate a regime with larger clusters. We use simulations of dynamic networks with different structures [58] to extract the effective parameters (mean cluster size $n\right>$, effective unit diffusivity $D_{n}$, steric radius $a_{n}$, contact volume $v_{c}$, and effective local fusion rate $k_{u}$) for the simplified social network model (details in SI Appendix).
When individual clusters are homogeneously filled upon each interaction ($\tau_{c}<\left\{\tau_{d},\tau_{\text{int}},\tau_{f}\right\}$), the solution in Eq. 3 can be used directly with the appropriate values of the parameters $a,v_{c},k_{u}$. However, for slowly diffusing particles or very large clusters ($\tau_{c}>\tau_{d}$), the source cluster may be only partly full of material. The average concentration in the source cluster, $h_{0}^{(d)}$, is then determined by the static solution (Eq. 1) within the cluster: $\ell_{0}h_{0}^{(d)}=(S_{d}(\sqrt{D_{p}/k_{d}})+1)/\langle n\rangle$. When another cluster fuses with the source cluster, fission may terminate the encounter before filling is complete ($\tau_{f}<\tau_{c}$). The amount transferred should then be scaled by the fraction of the new cluster that is filled prior to fission: $f=(S_{d}(\sqrt{D_{p}/k_{f}})+1)/\langle n\rangle$. These two corrections modify the dynamic Eq. 2 for the concentration in the contact zone and its solution $S$ as follows:
$\frac{dh_{c}(t)}{dt}$ $=k_{u}\left[\ell_{0}h_{0}^{(d)}-h_{c}(t)\right]f-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (4a)
$S$ $=S_{d}\left(\sqrt{D_{p}/k_{d}}\right)+\rho h_{c}(v_{c}+z),$ (4b)
where the first term in the updated $S$ accounts for material in the source cluster and the second term represents material in the rest of the network. Note that $h_{c}$ now depends on $h_{0}^{(d)},f$. When clusters are large and interactions are infrequent, the above solution reduces to the static network limit of Eq. 1. When the clusters are small, the source cluster is fully filled ($h_{0}^{(d)}\rightarrow h_{0}$) and so is each cluster that interacts with it ($f\rightarrow 1$). The general solution then approaches the fragmented social network limit of Eq. 3.
In Fig. 4, we show the spreading of material through simulated networks with increasing connectivity, spanning across the different regimes. In the limit of rapid decay (small $\tau_{\textrm{d}}$), material spreads within a single cluster of dimension $d$, with the amount delivered scaling as $S\sim\tau_{d}^{d/2}$ When particle diffusivity is fast and clusters are small, the source cluster fills faster than interactions can occur and the amount of material plateaus in the regime of $\tau_{c}<\tau_{d}<\tau_{\text{int}}$. For longer decay times, interaction between distinct clusters dominates the spread, and the total material scales linearly as $S\sim\tau_{d}$. When particle diffusion is slow (Fig. 4b), the cluster containing the source unit is able to offload material through interactions even before it is fully filled ($\tau_{\text{int}}<\tau_{c}$), and the plateau region narrows or disappears.
The simplified analytic model (Eq. 4) approximately matches simulations over a broad range of network connectivities and timescales (Fig. 4). However, some discrepancy occurs for networks that are close to the percolation transition (light green curve), particularly when the particle diffusivity is low. In this case, the approximation systematically overestimates material spreading through the network. This may be the result of transient fissions within a single cluster (which could temporarily hinder material spreading [22]), or steric inaccessibility of nodes buried within a cluster, neither of which are accounted for in our analytic models. Additionally, the percolation transition corresponds to a broad variability in cluster sizes, resulting in the source unit occasionally being trapped in very small clusters that limit material delivery.
The quantitative description of steady-state network filling (Eq. 4) can be generalized to other measures of material spread on a dynamic network. In particular, the decay time $\tau_{d}$ sets the timescale over which the spreading is assessed. The total amount of material delivered into the network over time is also well-approximated by the physical and social network models presented here (see Supplemental Fig. S1 in SI Appendix).
Discussion
The above calculations yield succinct, conceptually interpretable predictions for the rate of material spread in dynamic networks. We examine two limiting regimes: one of static, well-connected physical networks and one of socially-interacting homogeneous clusters. In physical networks, spreading increases with material diffusivity and network dimensionality. In social networks, material accumulates linearly in time and is limited either by the fusion rate between nearby clusters or the mobility of those clusters.
The models presented here are based on measurable structural and dynamic features. The static model requires network dimensionality, cluster size, and particle diffusivity, while the social model additionally needs the fusion rate and cluster mobility. Modern imaging techniques allow quantification of mitochondrial network structure and dynamics in a variety of cellular systems  [72, 42, 73, 14, 29]. Our results connect these morphological measurements to the dynamics of material spreading through the mitochondrial population, which is experimentally more challenging to assess. The scaling relationships could in principle be tested by quantifying the dispersion of locally photoconverted proteins through mitochondrial networks of different architectures [35].
This work brings us closer to understanding the functional implications of variable network connectivity for the spread of ions, lipids, proteins, and genetic information within mitochondria. Mammalian cells provide ample examples of mitochondrial network structures in different regimes, from fragmented networks in cells grown on excess glucose [35] or endowed with a Down’s syndrome genotype [34], to networks balanced on the cusp of percolation [14, 73, 58], to hyperfused networks observed in starved cells [35]. A quantitative framework for material spreading in different structures provides insight on the implications of these transitions for many biologically critical processes, such as the distribution of harmful reactive oxygen species [48], the ability to isolate or complement deleterious DNA mutations [74], and the tunneling of calcium ions, proton gradients, and ATP throughout the cell [75, 49, 50].
Our quantitative calculations focus on a simplified dynamic system, with a single source of material held at fixed concentration. Realistic biological scenarios are of course likely to involve additional complications, including interfering sources, non-trivial regulation at the source, or diffusive barriers within the mitochondria [62]. We also assumed that individual mitochondrial units are all identical in their fusion and fission behavior, whereas it is possible that some units remain isolated from the rest of the network, while others are more likely to engage in interactions. However, the minimal model presented here provides a basic building block from which more complicated scenarios can be constructed and which can serve as a null hypothesis for analyzing experimental observations.
The model makes general testable predictions that may be compared against future measurements. We predict that in a fragmented social network, spreading rates should be approximately the same for particles of different diffusivity. In highly-connected physical networks, branched architectures should give rise to more rapid material spread, with a steeper scaling in time, than linear networks. The models presented here could be extended to make quantitative predictions of network filling from multiple source points. For example, this approach could be used to explore how the mitochondrial network fills with calcium that enters at many localized contact sites with the endoplasmic reticulum [76]. Similarly, it enables predictions of how the distribution of mRNA localization and protein import sites on the mitochondrial surface [77, 78] translates into heterogeneity in protein levels within the mitochondria.
While we have focused on mitochondria, analogous problems of diffusive spreading arise for porous media in the geosciences [24], trophallactic networks in honeybees [4], and Alzheimers progression across connected regions in the brain [79]. By bringing together models of static physical networks and transiently interacting social networks, our results enable a quantitatively predictive link from network architecture to the rate of diffusive dispersion.
Acknowledgements.
Funding for this work was provided by NSF grant PHYS-2310229, by the Chan Zuckerberg Initiative, and by a grant from the UCSD Academic Senate. We are grateful to Nigel Goldenfeld and to the members of the CZI Theory Institute Without Walls for helpful discussions.
Supplemental Appendix
.1 Exact solution for static physical networks
For a stationary network structure, the total material content at steady-state can be computed analytically by solving the diffusion equation on each linear segment, then linking these solutions together at the junctions. The segment containing a source unit is separated into two segments on either side of the source. The diffusion equation and corresponding boundary conditions for each segment is:
$\begin{gathered}D_{p}\frac{d^{2}c_{n}(x)}{dx^{2}}-k_{d}c_{n}(x)=0,\\ c_{n_{1}^{i}}(x_{n_{1}^{i}})=c_{n_{2}^{i}}(x_{n_{2}^{i}})=c_{n_{3}^{i}}(x_{n_{3}^{i}}),\\ \sum_{j=1}^{\textrm{deg}(i)}\frac{dc_{n_{j}^{i}}}{dx}\bigg{\rvert}_{x_{n_{j}^{i}}}=0\end{gathered}$ (5)
where $n_{j}^{i}$ is the $j$-th edge connected to node $i$ and $x_{n_{j}^{i}}$ is the value of $x$ on that edge at node $i$. For each segment, $0<x<\ell_{n_{j}^{i}}$ and $x_{n_{j}^{i}}\in\left\{0,\ell_{n_{j}^{i}}\right\}$, where $\ell_{n_{j}^{i}}$ is the segment length. The middle equation defines continuity of the concentration fields, and the final equation conserves flux. This system has general solutions:
$c_{n}(x)=A_{n}\cosh(x/\lambda_{p})+B_{n}\sinh(x/\lambda_{p})$
allowing us to rewrite the boundary conditions as a large matrix equation $M\vec{\alpha}=\vec{v}$, where $\vec{\alpha}$ is a stacked vector of coefficients $A_{n},B_{n}$, and $\vec{v}$ is zero except for boundaries of the source unit. The total material in the network is found by integrating along all segments $m$ outside of the source unit, to yield $S=\sum_{m}\int_{0}^{\ell_{m}}c_{m}(x)\,\mathrm{d}x$.
Because the material delivered will depend on the choice of source unit within the network, we repeat this procedure for all possible choices of source unit and average the values of $S$. This result is then averaged over each replicate snapshot to give the exact solution (solid curves in Fig. 2).
.2 Continuum Approximation for Static Physical Networks
We model highly-connected, stationary, physical network structures as an effective $d$-dimensional sphere through which diffusive material spreads from a localized source.
.2.1 Selection of the effective domain size $R_{n}$
To extract the appropriate size of the continuum domain, we use the dimension $d$ and the mean graph distance between paris of nodes $\langle r_{g}\rangle$ in a given network. As we are approximating the network by a $d$-dimensional sphere, the mean distance between points in that sphere [80] should be equal to the mean graph distance. This yields the relationship:
$\begin{gathered}R_{n}=\frac{(2d+1)\langle r_{g}\rangle}{2d\beta_{d}},\\ \beta_{d}=\frac{2^{d+1}\Gamma(d+1)^{3}}{(d+1)\Gamma(\frac{d+1}{2})^{2}\Gamma(2d+1)}\end{gathered}$ (6)
However, for dimension less than 2, where diffusive search is compact and geometry-dependent [81], the position of the source within the domain becomes important.
For the linear case ($d=1$), specifically, a source point selected at random on a line of length $2R_{n}$, will on average be a distance $R_{n}/2$ from the nearest end. The steady-state distribution profile from this source can be solved analytically to give the total amount delivered:
$\begin{split}S=\frac{\lambda_{p}}{\ell_{0}}\left[\tanh(R_{n}/2\lambda_{p})+\tanh(3R_{n}/2\lambda_{p})\right]\end{split}$ (7)
An alternate approach is to solve for $S^{(1)}$: the material delivered from a central source in a linear domain of radius $R^{(1)}=R_{n}/2$, and separately $S^{(2)}$: material delivered from a central source in a domain of radius $R^{(2)}=2R_{n}-R_{n}/2$. Averaging these results together gives the exact expression in Eq. 7.
Analogously, for the $d$-dimensional case, a source point selected at random in a hypersphere of radius $R_{n}$ will be at a distance of $R^{(1)}=1/(d+1)R_{n}$ from the boundary. We compute $S^{(1)}$: the total material delivered from a central source in a hypersphere of radius $R^{(1)}$. Separately, we compute the result $S^{(2)}$ for a hypersphere of radius $(R^{(2)})^{d}=2(R_{n})^{d}-(R^{(1)})^{d}$. The average d-dimensional volume of the two hyperspheres is thus equal to the total volume. The two resulting solutions for delivered material are averaged together to give $S=(S^{(1)}+S^{(2)})/2$. Although this is not an exact solution for $d>1$, it provides a good approximation for material delivery in simulated networks, where source points are selected uniformly to start anywhere on the network (Fig. 2, dashed-square curves).
.2.2 Solving the continuum d-dimensional model (Eq. 1)
Starting with the general solution for the concentration [70]:
$c(r)=Ax^{\nu}I_{\nu}(x)+Bx^{\nu}K_{\nu}(x)$ (8)
we first apply the boundary conditions to solve for the constants $A,B$ as follows. The reflecting outer boundary sets:
$\begin{gathered}\left.\frac{dc}{dr}\right|_{R_{n}}=\frac{\sqrt{d}}{\lambda_{p}}\left.\frac{dc}{dx}\right|_{x_{2}}=Ax_{2}^{\nu}I_{\nu-1}(x_{2})-Bx_{2}^{\nu}K_{\nu-1}(x_{2})=0\\ \rightarrow\quad\frac{B}{A}=\frac{I_{\nu-1}(x_{2})}{K_{\nu-1}(x_{2})}\end{gathered}$ (9)
and the fixed-concentration inner boundary sets:
$\begin{gathered}c(a_{n})=c(x_{1})=C_{0}^{(d)}=A\left(x_{1}^{\nu}I_{\nu}(x_{1})+\frac{B}{A}x_{1}^{\nu}K_{\nu}(x_{1})\right)\\ \rightarrow\quad A=\frac{C_{0}^{(d)}x_{1}^{-\nu}}{I_{\nu}(x_{1})+\frac{B}{A}K_{\nu}(x_{1})},\end{gathered}$ (10)
where $C_{0}^{(d)}$ is related to the linear concentration of the fixed unit, $h_{0}$, as $M_{d}a_{n}^{d}C_{0}^{(d)}=\ell_{0}h_{0}=1$, and $M_{d}$ is the volume of a $d$-dimensional unit sphere (i.e. $\frac{4}{3}\pi$ in 3D). The total material delivered is then found through integration of the concentration field over space:
[…]
$=$ ${U_{d}}\left(\frac{\lambda_{p}}{\sqrt{d}}\right)^{d}A\left[-x_{1}^{1-\nu}\left(I_{\nu-1}(x_{1})-\frac{B}{A}K_{\nu-1}(x_{1})\right)\right]$
where $U_{d}$ is surface area of a $d$-dimensional unit sphere (i.e. $4\pi$ in 3D) and the $x_{2}$ term vanishes by the outer boundary condition. Substituting in the definitions of $A,C_{0}^{(d)}$, and $B/A$ we arrive at Eq. 1:
$S_{d}=$ $U_{d}\left(\frac{\lambda_{p}}{\sqrt{d}}\right)^{d}x_{1}^{1-2\nu}\left(\frac{1}{M_{d}a_{n}^{d}}\right)\frac{\left[-I_{\nu-1}(x_{1})+\frac{B}{A}K_{\nu-1}(x_{1})\right]}{\left[I_{\nu}(x_{1})+\frac{B}{A}K_{\nu}(x_{1})\right]}$ (12)
$=$ $\frac{d}{x_{1}}\frac{\left[-I_{\nu-1}(x_{1})+\frac{I_{\nu-1}(x_{2})}{K_{\nu-1}(x_{2})}K_{\nu-1}(x_{1})\right]}{\left[I_{\nu}(x_{1})+\frac{I_{\nu-1}(x_{2})}{K_{\nu-1}(x_{2})}K_{\nu}(x_{1})\right]}$
since $U_{d}/M_{d}$ is simply the dimension $d$ [82].
.3 Mean-field model for social networks
For social networks consisting of identical diffusing units that exchange contents upon interaction, we develop a continuum field-based model for the spatial spreading of material. This model can apply to small clusters as well as completely fragmented units, so long as the material distribution within each cluster is approximately homogeneous.
.3.1 Model for spatial material concentration (Eq. 2)
We derive the dynamic equation of a spatial mean concentration field, $h(\vec{r},t)$, defined as the average material concentration per unit at position $\vec{r}$ relative to the source cluster at time $t$. For uniform diffusing clusters in 3D, we assume spherical symmetry and express the concentration field as a function of the radial coordinate $r$. The material is carried by clusters with relative diffusivity $D$ and decays at rate $k_{d}$, giving a simple reaction-diffusion equation for positions outside the contact zone ($r>b$):
$\frac{dh(r,t)}{dt}=D\nabla^{2}h(r,t)-k_{d}h(r,t)$ (13)
Material exchange between non-source clusters has no effect on the average material concentration at the given position.
Two adjustments are made to Eq. 13 when clusters are within the contact volume $v_{c}$ ($a<r<b$), where they are able to undergo fusion with the source cluster. First, we assume that the contact region is very thin relative to the domain size, so that the concentration field can be regarded as a spatially uniform value: $h_{c}(t)$. Material leaves the contact zone via diffusive flux through its outer boundary. This flux also forms the boundary condition at $r=b$ for Eq. 13. The spatial density of material is $\rho h$, where $\rho$ is the density of mitochondrial units. The overall current leaking out of the zone can then be expressed as:
$I=-D\rho\left.\frac{dh(r,t)}{dr}\right|_{b}(4\pi b^{2})$ (14)
The per-unit loss of concentration is $I/(\rho v_{c})$, as we must divide the total escaping concentration by the number of units in the contact volume.
An additional adjustment accounts for the injection of material into the network upon fusion through the term $+k_{u}(h_{0}\ell_{0}-h_{c})$, where $k_{u}$ is the rate for each cluster to fuse with the source. This term arises from the fact that whenever a cluster fuses with the source cluster, both clusters leave the interaction with concentration $h_{0}=1/\ell_{0}$, meaning that the material per unit in the non-source cluster increases by ($h_{0}\ell_{0}-h_{c}$). With these modifications, we arrive at the dynamic equation for per-unit concentration within the contact zone:
$\frac{dh_{c}(t)}{dt}=k_{u}(h_{0}\ell_{0}-h_{c}(t))-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t)$ (15)
At steady-state, the time derivatives in Eq. 1315 are set to 0 to solve for the spatial profile $h(r)$, as described below.
.3.2 Selection of the steric and contact radii, $a,b$ and effective fusion rate, $k_{u}$
The continuum model requires defining inner and outer radii $a,b$ for interacting clusters. Since the inner radius $a$ represents a steric exclusion distance, we set $a=2R_{g}$, where $R_{g}$ is the average radius of gyration for clusters in the network. Rather than defining $b$ directly, we set the total volume $v_{c}$ available for fusion to a cluster by multiplying the average number of nodes available for fusion (those with degree below 3) by the contact volume per node. The contact volume per node describes the volume within which two individual nodes can undergo fusion; it is computed in detail in Ref. [58]. This procedure fixes $b$ through the relationship $v_{c}=\frac{4}{3}\pi(b^{3}-a^{3})$.
The effective fusion rate is also dependent on the number of nodes available for fusion. In our simulations, we have separate orientation-dependent rates for tip-tip (between two degree-1 nodes) and tip-side (between a degree-1 and degree-2 node) fusion: $k_{u1}(\Theta)$ and $k_{u2}(\Theta)$. The effective rate $k_{u}$ is calculated as a weighted sum of these two fusion types:
$k_{u}=$ $\langle k_{u1}(\Theta)\rangle\cdot x_{1}\left(\frac{x_{1}}{x_{1}+x_{2}}\right)+\langle k_{u2}(\Theta)\rangle\cdot\frac{2x_{1}x_{2}}{x_{1}+x_{2}}$ (16)
where the angle brackets denote averaging over orientations for a pair of nodes, given that they are close enough in space to fuse (within the contact volume per node) and $x_{1},x_{2}$ denote the average number of degree 1 and 2 nodes per cluster, respectively. Since all nodes in the incoming cluster have a chance to fuse with the source cluster while inside $v_{c}$, the tip-tip rate is multiplied by the number of degree-1 nodes in the incoming cluster and the probability to be near a degree-1 node from the source cluster, given a position inside $v_{c}$. The tip-side rate is weighted analogously, but contains two contributions. First, we include the number of degree-1 nodes in the incoming cluster multiplied by the probability to be near a degree-2 node from the source cluster. Second, we have the number of degree-2 nodes in the incoming cluster multiplied by the probability to be near a degree-1 node from the source cluster. Further details of the fusion and fission model and its dependence on the orientation and position of interacting units can be found in Ref. [58].
.3.3 Solving the mean-field model (Eq. 2,3)
We begin by writing down the general solution at steady-state, $h(r)=\frac{1}{r}\left(Ae^{r/\lambda}+Be^{-r/\lambda}\right)$ where $\lambda=\sqrt{D/k_{d}}$, with $D$ the relative cluster diffusivity. Applying the outer boundary condition gives:
$\begin{gathered}\left.\frac{dh}{dr}\right|_{R}=\frac{Ae^{R/\lambda}-Be^{-R/\lambda}}{R\lambda}-\frac{Ae^{R/\lambda}+Be^{-R/\lambda}}{R^{2}}=0\\ \rightarrow\quad\frac{B}{A}=e^{2R/\lambda}\left(\frac{R-\lambda}{R+\lambda}\right)\end{gathered}$ (17)
and the inner continuity condition gives:
$\begin{gathered}h(b)=\frac{A}{b}\left(e^{b/\lambda}+\frac{B}{A}e^{-b/\lambda}\right)=h_{c}\\ \rightarrow\quad A=\frac{(R+\lambda)e^{-b/\lambda}bh_{c}}{(R+\lambda)+(R-\lambda)e^{2(R-b)}}\end{gathered}$ (18)
Next we find the steady-state $h_{c}$ by setting $\frac{dh_{c}}{dt}=0$. First, solving for the current across the boundary gives:
$\frac{I}{k_{d}\rho v_{c}}=$ $-\frac{4\pi b^{2}\lambda^{2}}{v_{c}}\left.\frac{dh}{dr}\right|_{b}$ (19)
$=$ $-\frac{4\pi b^{2}\lambda^{2}}{v_{c}}A\left(\frac{e^{b/\lambda}-\frac{B}{A}e^{-b/\lambda}}{b\lambda}-\frac{e^{b/\lambda}+\frac{B}{A}e^{-b/\lambda}}{b^{2}}\right)$
$=$ $h_{c}z/v_{c}$
where the last step comes from plugging in the results for $A,\frac{B}{A}$ and we define:
$\frac{z}{4\pi b\lambda}=\frac{(\lambda-b)(R+\lambda)+(\lambda+b)(R-\lambda)e^{2(R-b)/\lambda}}{(R+\lambda)+(R-\lambda)e^{2(R-b)/\lambda}}$ (20)
Now solving for $h_{c}$ directly:
$\begin{gathered}k_{u}(1-h_{c})-I/\rho v_{c}-k_{d}h_{c}=0\\ h_{c}(1+k_{u}/k_{d}+z/v_{c})=k_{u}/k_{d}\\ h_{c}=\frac{k_{u}/k_{d}}{1+k_{u}/k_{d}+z/v_{c}}\end{gathered}$ (21)
The total concentration outside of the fixed cluster is then given by:
[…]
$=$ $\rho\left(h_{c}v_{c}+4\pi A\left[(r\lambda-\lambda^{2})e^{r/\lambda}-\frac{B}{A}(r\lambda+\lambda^{2})e^{-r/\lambda}\right]_{b}^{R}\right)$
$=$ $\rho h_{c}(v_{c}+z)$
as in Eq. 3. To get the total concentration including the fixed cluster, we add the the material per unit at the source ($h_{0}\ell_{0}=1$) multiplied by the size of the fixed cluster, $(\langle n\rangle-1)$ (ignoring the fixed unit itself). The limiting behaviors for short $\tau_{d}$, $\tau_{\text{enc}}$, and $\tau_{u}$ (indicated in Fig. 3) are each derived straightforwardly through expansion of $z$ for small $\lambda\ll R$.
.4 Time-dependent spread of material without decay
The results in the main text focus on the total material delivered to a network at steady-state, where such a state is reached through a balance of spreading and decay. Simulations and analytical approximations provide a comprehensive picture for this system (Fig. 4). The decay rate $k_{d}$ in this model plays the role of setting the timescale on which spreading is being considered.
To show the generality of the analytic approximations, we consider the related problem of time-dependent spreading without decay. With $k_{d}=0$, there is no steady-state but we can instead quantify the temporal evolution of material content in the network: $S(t)$. To probe this question, we run simulations using the same parameters as in Fig. 4 with $k_{d}=0$, tracking the total material delivered to the network over time. Fig. S1 shows that the combined steady-state analytical solution (Eq. 4), provides a good approximation to the time-dependent spreading $S(t)$ when evaluated using the corresponding effective decay rate: $k_{d}=1/t$.
Figure S1: The total material in the network plotted as a function of time for a system with no material decay, with $S(t=0)=0$. Solid lines correspond to simulated networks, filling with material over time. Simulation parameters are identical to Fig. 4 but with $k_{d}=0$. Dashed curves give the combined analytic steady-state solution for physical and social networks, as in Fig. 4, plotted with $t=\tau_{d}$ as the effective spreading time.
.5 Simulation Methods
.5.1 Dynamic network simulations
We start with the dynamic network simulation framework described in [58]. Edge-units are spherocylinders with length $\ell_{0}=0.5$ and have steric exclusion radius $r_{s}=0.15$. All $N_{0}=250$ units are confined by a reflecting domain boundary of radius $R=5$. To generate the network structures, we use variable tip-tip and tip-side fusion rate constants $k_{u1},k_{u2}$ as indicated in the text, with the fission rate constant set to $k_{f}=1$ to normalize time units throughout. Fusion between two nodes can occur when the nodes are separated by less than $2r_{c}$, where $r_{c}=0.2$ is the contact radius. Random motion of units and clusters is realized by a diffusivity of $1$ applied at each node, giving $D_{1}\approx 0.5$ per isolated unit. Assuming a fission rate on the order of $k_{f}=1/(2\text{min})$, this dimensionless diffusivity corresponds to approximately $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$.
Diffusion of material through the network is accomplished using finite-volume simulations built on top of the dynamic network simulation. Material concentration fields are discretized on a per-unit basis, with the change in concentration across each junction (node) in the network calculated as:
$\frac{dc_{n_{j}^{i}}}{dt}=\frac{D_{p}}{\ell_{0}^{2}}\sum_{k=1}^{\textrm{deg}(i)}(c_{n_{k}^{i}}-c_{n_{j}^{i}})$ (23)
where $D_{p}=4800$ is the default particle diffusivity and $c_{n_{j}^{i}}$ is the per-unit concentration on the $j$-th edge connected to node $i$. Material exchange is thus only possible across connected (fused) units. Decay is applied uniformly across the network as $dc/dt=-k_{d}c$. The simulations proceed through forward Euler stepping with a timestep of $\Delta t=10^{-4}$. After allowing the network structure to equilibrate for $2\times 10^{6}$ simulation steps, we initialize the concentration of all units at $c=0$, except for one source unit chosen at random, whose per-unit concentration is fixed to $c=\ell_{0}h_{0}=1$. We then run each spreading simulation for at least 5 times the decay time ($5\tau_{d}$), averaging the total material over the second half of the simulation. We further average over 9 choices of the source node per simulation and at least 3 independent simulation replicates.
.5.2 Interacting spheres simulations for social networks
For simulations of the simplified model for highly fragmented mitochondria, we take an agent-based approach with $N_{0}=251$ identical spherical units of steric exclusion radius $a/2=0.15$. These spherical units are initially distributed uniformly inside a reflecting spherical boundary of radius $R=5$ and are allowed to diffuse with diffusivity $D_{1}=0.5$. One unit is selected as the source and is assigned constant concentration $c=1$. Transient fusion between two units occurs at rate $k_{u}$ whenever the centers of the spheres are separated by less than $b$, where $b=0.4$ represents the contact distance. Upon fusion, the sphere concentrations $c_{1},c_{2}$ are updated to $c_{1}^{\prime}=c_{2}^{\prime}=(c_{1}+c_{2})/2$ to represent equilibration. Fusions are immediately followed by fission. If a non-source unit fuses with the source unit, its concentration is raised to $c=1$. Decay is applied uniformly across the non-source units with $c(t+\Delta t)=c(t)e^{-\Delta t/\tau_{d}}$ where the timestep $\Delta t=10^{-2}\cdot\min(10^{-1},k_{u}^{-1},\tau_{d})$. We run each simulation until the simulation time is at least $t=50$ for $\tau_{d}\leq 1$ and at least $t=2000$ for $\tau_{d}>1$, and average the total material over the last quarter of the simulation.